OpenStudy (anonymous):

What's 3 3/4 x (-1 1/2)? I keep on getting the wrong answer

OpenStudy (mathmale):

Please convert 3 3/4 into an improper fraction. Then, separately, please convert (-1 1/2) into an improper fraction. Share your results here.

OpenStudy (anonymous):

15/4

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

mathmale?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@Yuna123

OpenStudy (mathmale):

Nice work. Let's now multiply those improper fractions together:\[\frac{ 15 }{ 4 }*\frac{ -3 }{ 2 }\]

OpenStudy (anonymous):

-12/8?

OpenStudy (anonymous):

or 12/8

OpenStudy (mathmale):

\[(\frac{ 15 }{ 4 })(\frac{ -3 }{ 2 })\] Unfortunately, nothing here can be cancelled/ reduced.

OpenStudy (mathmale):

Try multiplying these together. What's 15 times -3?

OpenStudy (anonymous):

-12

OpenStudy (mathmale):

15 times 2 is 30; 15 times 4 is 60. Try again. What's 15 times 3? what's 15 times -3?

OpenStudy (anonymous):

oh wait is the the rule for + and -? sub and add?

OpenStudy (anonymous):

so -45?

OpenStudy (mathmale):

You are multiplying here. (-1)(-1)=1 (-1)(1)=-1 Yes, -45. Now multiply the 2 denominators together. 4(2)=?

OpenStudy (anonymous):

8

OpenStudy (mathmale):

Now write your final result, remembering that you were dealing with fractions and multiplying two improper fractions together.

OpenStudy (anonymous):

-6/1?

OpenStudy (mathmale):

How did you obtain that? You had a numerator of -45 and a denominator of 8.

OpenStudy (whpalmer4):

@loveIt this is just straightforward multiplication. \[(\frac{15}4)(\frac{-3}2) = \frac{(15)(-3)}{(4)(2)} = \frac{-45}{8} = -\frac{45}8\]

OpenStudy (whpalmer4):

if we factor the numerator and denominator, we can see there are no common factors, so this cannot be simplified further: \[-\frac{45}8=-\frac{3*3*5}{2*2*2}\] we can make this into an improper fraction, however. \[45/8 = 5 \text{ with remainder of }5\]so \[-\frac{45}8=-5\,\frac{5}8\]